# High-precision measurements of electric-dipole-transition amplitudes in excited states of 208Pb using Faraday rotation spectroscopy

### v1

@@ V Discussion
After including statistical and systematic errors, and combining our Faraday and transmission results as described in the previous section, we arrive at the following final experimental values for the two Pb matrix elements studied here -- for the $(6p^{2})^{3}P_{1}\to(6p7s)^{3}P_{0}$ 368 nm transition we find: $\langle||E1||\rangle_{368}=1.897(11)$ a.u., and for the $(6p^{2})^{3}P_{2}\to(6p7s)^{3}P_{1}$ 406 nm transiton we find: $\langle||E1||\rangle_{406}=3.008(22)$ a.u.. From Table 1 it is clear that the temperature uncertainty of our vapor cell sample dominates the final combined error in each case. As it turns out, for any given temperature error, we find a three-fold smaller fractional uncertainty contribution if we consider the ratio of our E1 matrix element measurements (there is partial, but not complete cancellation of Boltzmann factors and associated uncertainties). However, this significant reduction in the overall error budget is counterbalanced by the increased error resulting from taking the ratio of our two otherwise uncorrelated measurements. In the end, we find that
@@
and we see that the fractional error in this quantity is roughly the same as that for our individual matrix element determinations.
We are now in a position to compare our new measurements to previous experimental work as well as the latest theoretical work. Substantial progress has been made in recent years in ab initio modeling of electron wavefunctions in multi-valence atomic systems. Current, state-of-the-art ‘CI+all order’ computational techniques have yielded both improved precision, and excellent agreement with recent benchmark experiments in three-valence indium and four-valence lead from our group Vilas2018 ; Maser2019 ; Safronova2025 . We here quote the following recommended ab initio theory values for our two matrix elements. For the 368 nm transition: $\langle||E1||\rangle_{th}=1.92(6)$ a.u. Maser2019 and for the 406 nm transition: $\langle||E1||\rangle_{th}=2.99(9)$ a.u. SafronovaPC . The estimated 3% theory uncertainties arise from comparison of results using an alternative CI-based approach, as well as estimates of the size of several small uncalculated corrections SafronovaPC . Our results now establish an important new experimental sub-1%-precision benchmark against which to compare future theory work. We note that in principle, given parallel theoretical treatments of these two matrix element calculations, a theoretical ratio of these quantities (in analogy to Eq. 10) will exhibit a somewhat reduced fractional error since both transitions are associated with common fine structure manifolds.
@@ VI Concluding Remarks
We have completed measurements of the $(6p^{2})^{3}P_{1}\to(6p7s)^{3}P_{0}$ and $(6p^{2})^{3}P_{2}\to(6p7s)^{3}P_{1}$ excited-state transitions in 208Pb offering significant improvement in precision compared with previous experimental measurements. Our measured results show excellent agreement with the ab initio theory, and our ratio of E1 values provides a sub-1 % experimental benchmark for theory. The described method of measuring excited-state transition amplitudes using high-precision Faraday and transmission spectroscopy opens up the possibility for a range of future measurements in other heavy element species. As a notable example, the low-lying level structure of thallium, which includes an accurately known ground-state infrared M1 transition, points to a similar scheme of E1/M1 comparison through Faraday rotation measurements of multiple low-lying thallium excited states.
Currently we are also pursuing additional experimental benchmarks through measurements of lead polarizabilities using transverse atomic beam spectroscopy in a similar fashion to our earlier work in thallium and indium Ranjit2013 ; Augenbraun2016 ; Vilas2018 ; Doret2002 . In the course of this atomic-beam-based work, we intend to pursue transverse laser cooling of a Pb atomic beam making use of the (optical cycling) 368 nm $(6p^{2})^{3}P_{1}\to(6p7s)^{3}P_{0}$ transition that has been studied in the present work. This first demonstration of lead laser cooling would represent a key initial step to forming a sample of laser-cooled lead atoms suitable ultimately for assembly of cold Pb-based diatomic molecules for potential next-generation EDM studies.
@@ Appendix A A: Reduced matrix elements
In all cases, our lasers are linearly polarized perpendicular to the quantization axis defined by the propagation direction and the DC Faraday magnetic field. Thus we can parametrize our transitions as due to equal parts $\sigma_{\pm}$ circular polarization components. We label our matrix elements as $|n,J,m\rangle$, where $n$ refers to the atomic state, $J$ to the total electronic angular momentum, and $m$ to its projection. We note that where there are multiple initial states involved we assume that (such as for the $(6s^{2})^{3}P_{1}$ and $(6p^{2})^{3}P_{1}$ excited states relevant for the 368 nm and 406 nm transitions respectively) all sub-levels are equally populated according to the relevant thermal Boltzmann factor.
@@ Appendix B B: Magnetic field measurement
By symmetry, the second lock-in peak corresponds to a $\sigma_{-}$ M1 transition followed by a $\sigma_{+}$ E1 transition that is shifted by $\delta f_{368}^{\sigma_{-}}=-\delta f_{368}^{\sigma_{+}}$. Ultimately, the observed frequency splitting is conveniently ‘amplified’ more than fourfold without loss of accuracy via this two-step technique and can be expressed in terms of the Landé $g$-factor, $g_{J}$, the Bohr magneton, the applied field, and Planck’s constant as
@@ Appendix C C: Line shape fitting functions
where the leading terms are $a_{0}=\sqrt{\pi}e^{-x^{2}}$, $a_{1}=4xF(x)-2$, $a_{2}=\sqrt{\pi}e^{-x^{2}}(1-2x^{2})$, $a_{3}=-\frac{1}{6}(8-24xF(x)-8x^{2}+16x^{3}F(x))$

### v2

@@ V Discussion
After including statistical and systematic errors, and combining our Faraday and transmission results as described in the previous section, we arrive at the following final experimental values for the two Pb matrix elements studied here -- for the $(6p^{2})^{3}P_{1}\to(6p7s)^{3}P_{0}$ 368 nm transition we find: $\langle||E1||\rangle_{368}=1.897(11)$ a.u., and for the $(6p^{2})^{3}P_{2}\to(6p7s)^{3}P_{1}$ 406 nm transiton we find: $\langle||E1||\rangle_{406}=3.008(22)$ a.u.. From Table 1 it is clear that the temperature uncertainty of our vapor cell sample dominates the final combined error in each case. As it turns out, for any given temperature error, we find a three-fold smaller fractional uncertainty contribution if we consider the ratio of our E1 matrix element measurements (there is partial, but not complete, cancellation of Boltzmann factors and associated uncertainties). However, this significant reduction in the overall error budget is counterbalanced by the increased error resulting from taking the ratio of our two otherwise uncorrelated measurements. In the end, we find that
@@
and we see that the fractional error in this quantity is roughly the same as that for our individual matrix element determinations. Notably, the reduced dependence on cell temperature in the ratio value offers a useful cross-check on our temperature determination procedure generally. A significant systematic temperature error beyond our quoted temperature uncertainties would have biased both E1 amplitude measurements in a common-mode fashion, but would not be apparent in the ratio value where the temperature-related error contribution is small. The fact that both the individual measurements and the ratio are in very good agreement with the theory predictions reinforces our confidence in the temperature determination and quoted uncertainty.
We are now in a position to compare our new measurements to previous experimental work as well as the latest theoretical work. Substantial progress has been made in recent years in ab initio modeling of electron wavefunctions in multi-valence atomic systems. Current, state-of-the-art ‘CI+all order’ computational techniques have yielded both improved precision, and excellent agreement with recent benchmark experiments in three-valence indium and four-valence lead from our group Vilas2018 ; Maser2019 ; Safronova2025 . We here quote the following recommended ab initio theory values for our two matrix elements. For the 368 nm transition: $\langle||E1||\rangle_{th}=1.92(6)$ a.u. Maser2019 and for the 406 nm transition: $\langle||E1||\rangle_{th}=2.99(9)$ a.u. SafronovaPC . The estimated 3% theory uncertainties arise from comparison of results using an alternative CI-based approach, as well as estimates of the size of several small uncalculated corrections SafronovaPC . Our results now establish an important new experimental sub-1%-precision benchmark against which to compare future theory work. We note that in principle, given parallel theoretical treatments of these two matrix element calculations, a theoretical ratio of these quantities (which, in analogy to Eq.10, would produce the ratio: 1.56) will exhibit a somewhat reduced fractional error since both the lower and upper levels of these transitions are associated with common fine structure manifolds.
@@ VI Concluding Remarks
We have completed measurements of the $(6p^{2})^{3}P_{1}\to(6p7s)^{3}P_{0}$ and $(6p^{2})^{3}P_{2}\to(6p7s)^{3}P_{1}$ excited-state transitions in 208Pb offering significant improvement in precision compared with previous experimental measurements. Our measured results show excellent agreement with the ab initio theory, and our ratio of E1 values provides an experimental benchmark for theory with sub-1% precision. Our experimental method of measuring excited-state transition amplitudes using high-precision Faraday and transmission spectroscopy of E1 and M1 transitions opens up the possibility for future measurements in other heavy element species. As a notable example, the low-lying level structure of thallium, which includes an accurately known ground-state infrared M1 transition, points to a similar scheme of E1/M1 comparison through Faraday rotation measurements of multiple low-lying thallium excited states with the expectation of similar sub-1% precision.
Currently we are also pursuing additional experimental benchmarks through measurements of lead polarizabilities using transverse atomic beam spectroscopy in a similar fashion to our earlier work in thallium and indium Ranjit2013 ; Augenbraun2016 ; Vilas2018 ; Doret2002 . In the course of this atomic-beam-based work, we intend to pursue transverse laser cooling of a Pb atomic beam making use of the (optical cycling) 368 nm $(6p^{2})^{3}P_{1}\to(6p7s)^{3}P_{0}$ transition that has been studied in the present work. The laser cooling of lead would be a first step towards the assembly of cold Pb-based diatomic molecules for potential next-generation EDM studies.
@@ Appendix A A: Reduced matrix elements
In all cases, our lasers are linearly polarized perpendicular to the quantization axis defined by the propagation direction and the DC Faraday magnetic field. Thus we can parametrize our transitions as due to equal parts $\sigma_{\pm}$ circular polarization components. We label our matrix elements as $|n,J,m\rangle$, where $n$ refers to the atomic state, $J$ to the total electronic angular momentum, and $m$ to its projection. We note that where there are multiple initial states involved we assume that (such as for the $(6s^{2})^{3}P_{1}$ and $(6p^{2})^{3}P_{2}$ excited states relevant for the 368 nm and 406 nm transitions respectively) all sub-levels are equally populated according to the relevant thermal Boltzmann factor.
@@ Appendix B B: Magnetic field measurement
By symmetry, the second lock-in peak corresponds to a $\sigma_{-}$ M1 transition followed by a $\sigma_{+}$ E1 transition that is shifted by $\delta f_{368}^{\sigma_{-}}=-\delta f_{368}^{\sigma_{+}}$. Ultimately, the observed frequency splitting is conveniently ‘amplified’ more than fourfold without loss of accuracy via this two-step technique and can be expressed in terms of the Landé $g$-factor, $g_{J}$, the Bohr magneton, $\mu_{\mathrm{B}}$, the applied field, $B$, and Planck’s constant, $h$, as
@@ Appendix C C: Line shape fitting functions
where the leading terms are $a_{0}=\sqrt{\pi}e^{-x^{2}}$, $a_{1}=4xF(x)-2$, $a_{2}=\sqrt{\pi}e^{-x^{2}}(1-2x^{2})$, $a_{3}=-\frac{1}{6}(8-24xF(x)-8x^{2}+16x^{3}F(x))$,…